Grade 9Mechanics

IN A TYPICAL BUGGI (A WOODEN PLATE IS FIXED ON THE REAR ON WHICH ONE PERSON CAN SIT) A BUGGI OF MASS 200KG IS MOVING A SPEED OF 10KM/H AS IT OVERTAKES A SCHOOL BOY WALKING AT A SPEED OG 4KM/H .THE BOY SITS ON THE WOODEN PLATE,IF THE MASS OF THE BOY IS 25 KG THE NEW VELOCITY OF THE BUGGI IS...()
A).28/3KM/H B).28/3M/S C).14/3KM/H D).14/3M/S

sha fisal
8 Years agoGrade 9

1 Answer

Surodeep Sheth
8 Years ago
As linear momentum is conserved,
m1 u1 + m2 u2 = (m1 + m2)v
200*10 + 25*4 = (200+25) v
2100/225 = v
v=28/3 km/h